OpenStudy (anonymous):

Help Pleaseee!(: A particle's position coordinates (x, y) are (3.0 m, 5.0 m) at t = 0; (6.0 m, 9.0 m) at t = 2.0 s; and (15.0 m, 15.0 m) at t = 5.0 s. ==>Find the magnitude of the average velocity from t = 0 to t = 5 s.

OpenStudy (tkhunny):

What's stopping you from using the distance formula? That should provide magnitude. Distance = Rate * Time When was the last time you pulled that formula out of your hat?

OpenStudy (anonymous):

i tried it and i got the wrong answer. I don't know what I'm doing wrong. @tkhunny

OpenStudy (tkhunny):

Show me what you did and what happened.

OpenStudy (anonymous):

\[vector:V(sub,m)= \frac{ r(sub,f) - r(sub,i) }{ \Delta t}\] \[=\frac{ (15,15)-(3,5) }{ 2 } =(1.5,2.0)m/s\] \[V(sub,m)=\sqrt{5^2+6^2} = 7.8 m/s\] @tkhunny

OpenStudy (tkhunny):

I don't see a distance formula in there. Thus, your equations still have no idea how far we have traveled. The notation is a little funny, but we should get \((15,15) - (3,5) = \sqrt{(15-3)^{2}+(15-5)^{2}}\). I can't really tell what your notation is trying to tell me.

OpenStudy (anonymous):

Oh I did it wrong I guess, I tried to follow an example.

OpenStudy (tkhunny):

Too bad the example apparently wasn't clear enough. You understand the use of the distance formula here?

OpenStudy (anonymous):

yes... could you guide me?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

The arithmetic is already spelled out up above. Please do the arithmetic indicated. What's the square root of 244?

OpenStudy (anonymous):

15.62? but that answer was wrong as well.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Please learn a little patience and stop looking at the answers. Let's solve the problem and THEN look at the answers. Moving from (3,5) to (15,15) is a magnitude of \(\sqrt{(15-3)^{2} + (15-5)^{2}} = \sqrt{12^{2} + 10^{2}} = \sqrt{144 + 100} = \sqrt{244} = 2\sqrt{61} = 15.6\). This is the total magnitude of the change in position. Since this is not what the problem statement wants, we are not yet done. Using the old, familiar forumla, Distance = Rate * Time, and solving for Rate, we get Rate = Distance / Time. We then compute the average velocity like that: \(Average\;Velocity = \dfrac{Total\;Displacement}{Total\;Time}\). Once we have this value, we should be done.

OpenStudy (anonymous):

so it would be 15.6/5?

OpenStudy (tkhunny):

Yup.

OpenStudy (anonymous):

Thank you so much!!!(: @tkhunny

OpenStudy (tkhunny):

We got it? Excellent. Now I 'm really scratching my head about that example you tried to follow. Oh, well. Good luck with that. :-)

OpenStudy (anonymous):

Yeah, it was an example my friend had... Thank you though! I really appreciate the help! @tkhunny